OpenStudy (anonymous):

The length of the base of a square pyramid is 15 inches. The slant height is 10 inches. What is the surface area of the pyramid?

OpenStudy (anonymous):

@eliassaab

hartnn (hartnn):

there is a direct formula, can you use it or you need to derive the formula ?

hartnn (hartnn):

\(\large S A = a^2+2a \sqrt {\dfrac{a^2}{4}+h^2}\)

hartnn (hartnn):

where a = base length = 15 h = height

hartnn (hartnn):

oh, since slant height is directly given, we can directly take \(\Large SL = \sqrt {\dfrac{a^2}{4}+h^2}=10 \) and your formula simplifies to \(\Large SA =a^2+2a\times SL\) a = 15, SL = 10 plug in! :)